OpenStudy (wolf1728):

18/3/3 - How to evaluate that? You can say 18/3 is 6 and dividing that by 3 = 2. OR You could say 3/3 = 1 and dividing 18 by that = 18. How should that be evaluated?

OpenStudy (ikram002p):

priority goes to the first operation :D

ganeshie8 (ganeshie8):

`left to right` precedence

ganeshie8 (ganeshie8):

a/b/c = (a/b)/c according to `left to right` precedence rule

OpenStudy (ikram002p):

yep but its priority \ \ have equal priority thus the first one have the right to be executed first xD

ganeshie8 (ganeshie8):

when you have equal priority, execute it from `left to right`

OpenStudy (wolf1728):

 So, if it is written like this then it is top to bottom priority?

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=18%2F3%2F3%2F1

ganeshie8 (ganeshie8):

looks like \[\large \dfrac{\dfrac{18}{3}}{3} = \dfrac{\left(\dfrac{18}{3}\right)}{3}\]

OpenStudy (wolf1728):

Okay then how about simplifying an algebraic expression like this? a*x / y / y Does it get written as (a*x / y) / y ?

ganeshie8 (ganeshie8):

multiplication and division have same precedence in PEMDAS

ganeshie8 (ganeshie8):

so we just execute everything from left to right

ganeshie8 (ganeshie8):

a*x / y / y = ((a*x)/y)/y

ganeshie8 (ganeshie8):

first operation would be multiplication

ganeshie8 (ganeshie8):

because it is on the left most

OpenStudy (wolf1728):

Isn't there a way to simplify that better? I always hated working with "3 level" fractions (or whatever the heck they might be called.

ganeshie8 (ganeshie8):

simplify using fraction properties ?

ganeshie8 (ganeshie8):

\[\large \dfrac{\dfrac{a*x}{y}}{y} = \dfrac{\left(\dfrac{ax}{y}\right)}{y}\] dividing \(y\) top and bottom gives \[\large \dfrac{ax}{y^2} \]

OpenStudy (wolf1728):

This was typed before your last response: ************************************************************************ Yes. I mean if an algebraic formula had to be written ((a*x) / y) / y Isn't there a way to make it look a little better? ************************************************************** Okay so ((a*x) / y) / y can be written much neater as: (a*x) / y^2 That's a LOT better.

ganeshie8 (ganeshie8):

yes that definitely looks better

OpenStudy (wolf1728):

Sure does. I hate those "3 level" fractions. Is there a specific algebra term for those?

ganeshie8 (ganeshie8):

i think they call them `complex fractions` in shcool textbooks

OpenStudy (wolf1728):

I remember way back when - I'd always get those messed up. (Seems they still present problems for me).

ganeshie8 (ganeshie8):

lol yeah i had to rely on wolfram to figure out the precedence

OpenStudy (wolf1728):

Okay - well thanks again.

ganeshie8 (ganeshie8):

np :) it is better to use parenthesis always to avoid the possible ambiguity

OpenStudy (wolf1728):

Definitely. Well, time to close the question. :-)

ganeshie8 (ganeshie8):

\[\large \dfrac{\dfrac{18}{3}}{3} = \dfrac{\left(\dfrac{18}{3}\right)}{3}\] i prefer the right side expression eventhough both equivalent

ganeshie8 (ganeshie8):

:)